# Compare without Despair: Reliable Preference Evaluation with Generation separability

### v2

@@ 1 Introduction
Finally, as a direct application of separability, we extend it to ELO, a rating system based on pairwise comparisons, now used widely for LLM generations (Chiang et al., 2024). By modifying the ELO rating update function to account for the separability of each new instance, our separability-ELO ratings provide more nuanced comparisons (§4). Overall, separability offers a reliable signal in the noisy landscape of generative evaluation via pairwise ratings, provides insights into benchmarks and test instances, and complements existing ranking measures for robust preference evaluation. Our code and data will be publicly released. [^1] https://github.com/dill-lab/separability
@@
At a very high level, there are four common scenarios which may occur in comparing generations from two models, which we highlight in Figure 2. Scenarios 1, 2 and 3 all depict output sets where any sample from model $A$ is expected to be very distant from any sample from model $B$ (i.e. low cross-alignment). It is easy to distinguish the two models under scenarios 1 and 3—if two generations are very different, they must be from different models. Under scenario 2, generations from the same model are also far from each other (i.e. the self-alignment is also low), which makes the overall output sets hard to distinguish. In contrast, scenario 4 depicts a situation where both self- and cross-alignments are high; all generations, regardless of the model they came from, are similar, making it hard for the rater to distinguish the models’ output sets.
@@
Figure 3: Histograms of separability distributions for summarization (Left) and translation (Right). For similar model pairs, CNN/DailyMail for news summarization and translation from a high-resource language (German) have lower average separability compared to SAMSum for dialogue summarization and translation from a lower-resource language (Czech). We use length-adjusted BERTScore Zhang et al. (2019) as the similarity metric (defined in Appendix A) for summarization and BLEU Papineni et al. (2002) for translation.
@@ 2.1 Calculating Generation Alignments
where $s:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a text similarity metric such as ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), or BLEU (Papineni et al., 2002). Intuitively, the alignment score $\mathcal{A}_{A,B}^{i}$ measures the expected similarity between an output from $m_{A}$ and an output from $m_{B}$, parameterized by $s$. A high value of $\mathcal{A}_{A,B}^{i}$ indicates high similarity (low variability) between the generations of the two models. Different similarity metrics can be used for different tasks. In cases where a user cares about fine-grained lexical differences, a metric such as BLEU may be suitable. On the other hand, if fine-grained lexical differences are less important than coarser semantic differences, metrics such as BERTScore or word embedding cosine similarity would be more suitable. We use a variation of BERTScore with a length-adjustment (defined in Appendix A), unless otherwise noted.
@@ 2.2 Calculating Generation separability
separability can take values in $[-1,1]$. [^4] We apply min-max normalization to constrain the range of alignments to $[0,1]$. In practice, however, cross-alignment is usually lower than self-alignment, making $\delta_{A,B}^{i}\in[0,1]$.
Figure 4: separability distributions for ART and BiSECT. We use length-adjusted BERTScore here (defined in Appendix A) as the similarity metric. separability has higher variance, especially for BiSECT, largely caused by differences in instruction prompt interpretation; see Appendix C.
@@ 2.3 Computing separability on Generation Benchmarks
We compute separability for various generation tasks under 6 different benchmarks using 3 model pairs, and demonstrate how separability can allow model developers and users to visualize and understand how much a model pair’s generations differ on a particular dataset. Figure 3 shows the empirical separability distributions on two summarization benchmarks (left): CNN/DailyMail (Nallapati et al., 2016) and SAMSum (Gliwa et al., 2019), and two machine translation benchmarks (right): Czech to English and German to English from the WMT-19 dataset (Barrault et al., 2019). Figure 4 shows the empirical separability distributions for abductive reasoning and sentence simplification; we use the ART (Bhagavatula et al., 2019) and BiSECT (Kim et al., 2021) benchmarks respectively. In each case, we compare three model pairs: GPT-3.5 vs. Vicuna-7B (LMSys, 2023), Vicuna-7B vs. Mistral-7B (Jiang et al., 2023), and GPT-3.5 vs. FLAN-T5-XXL (Longpre et al., 2023). Appendix A contains additional details about the generation settings. In Appendix C, we present examples of low and high separability generations corresponding to each of these tasks.
@@
We conduct a human study to verify our formulation for separability as a meta-evaluation measure of preference rating consistency for a given instance $\mathbf{x}^{i}$ and a pair of generative models, $m_{A}$ and $m_{B}$. Given that generation sets corresponding to low separability instances are harder to distinguish, we hypothesize that preference judgments from raters on those sets will be inconsistent. In other words, raters will not consistently prefer the same model’s generation for any pair of generations sampled from low separability instances.
Figure 6: Proportion of rating sets with each value in the range of consistency corresponding to different separability ranges. Ratings are not aggregated across annotators for a test instance here. For each model pair and dataset configuration, the support is divided into equal sized ranges. The proportion of perfectly consistent ratings increases, and the proportion of inconsistent ratings decreases in higher separability ranges.
@@ 3.1 Rating Consistency
We define consistency of preference judgments as the average ratings from raters over $N$ sampled pairs. For an input instance $\mathbf{x}^{i}$, we sample $N$ generations $\mathbf{y}^{i}\in\mathcal{Y}$ each from models $A$ and $B$ to obtain a set of paired generations $\mathcal{P}_{AB}(\mathbf{x}^{i})=\left\{\left(\mathbf{y}^{i,j}_{A},\mathbf{y}^{i,j}_{B}\right)\right\}_{j=1}^{N}$.
@@ 3.2 Study Protocol and Settings
We conducted a human study with raters hired from Amazon Mechanical Turk. Each human intelligence task (HIT) consisted of reading a source text (in our case, a news article or a dialogue) and $N=5$ pairs of generated summaries. [^5] While we performed our experiments with summarization, we expect our results to hold for other tasks as well. In addition, the annotators are performing $5$ cross-model comparisons as opposed to $25$ when calculating separability, but we find that $5$ comparisons suffice. For each summary pair, raters were asked to select which summary they preferred, with the option of picking no preference. The HIT interface and further details can be found in Appendix B.
@@
Figure 8: Mean consistency of human and auto-rater preference judgments increases with separability. Mean consistency is computed over all $600$ HITs collected. Consistency for a particular test instance is aggregated over annotators by taking the mean of each individual annotator’s rating consistency.
@@ 3.4 Higher separability Instances Receive Consistent Auto-Rater Ratings
LLM-based automatic raters have been rising in popularity (Chang et al., 2024) and are being used to replace human raters in many preference evaluation setups. We ask: do auto-raters produce similar patterns of consistency as humans when making preference judgments? We repeat our experiments using the same 600 instances in §3.3 with auto-raters provided by AlpacaEval (Dubois et al., 2024) Each test instance is judged by three auto-raters, which have the highest correlations with humans. [^6] https://github.com/tatsu-lab/alpaca_eval. Since these raters cannot give tie judgments, the only possible consistencies are 00 or $1$; other details are in Appendix A.
@@ 4 Applying separability to ELO
We compute ELO and separability-weighted ELO (sep-ELO; Equation 6) using data from our 600 human evaluation HITs as preference judgments (§3). To calculate these ratings, we sample one rating for each input from our pool of ratings. We compute confidence intervals with the bootstrap method with $100$ trials. Figure 9 shows that sep-ELO has narrower gaps in model ranking, suggesting that models are more similar under adjustments to consistency of judgments (or, separability). We acknowledge that our results are a pilot due to the limited number of ratings we could use in our computation. However, we expect sep-ELO can reveal reliable trends even when applied to larger sets of preference data such as LMSYS [^9] https://lmsys.org/.
@@
Beyond sep-ELO, separability values could be used for adversarially filtering test sets Bras et al. (2020). Not only would this lead to fine-grained comparisons between models, but could also lead to obtaining cost- and time-efficient human ratings. However, some caution is to be urged since such filtering may lead to biases Schwartz and Stanovsky (2022), since low separability instances can still contain valuable information. Instead, we recommend importance weighting instances by separability when sampling instances for human judgments, in a similar manner as it is used in ELO ratings. Alternatively, a stratified sampling approach from different separability ranges could ensure a more robust preference evaluation scheme.
@@ Acknowledgments
We would like to thank several members of USC NLP for their valuable feedback, insights, and help in designing and testing our human evaluation: Jaspreet Ranjit, Risha Surana, Xinyue Cui, Yoonsoo Nam, Keyu He, Joseph Liu, Matthew Finlayson, Brihi Joshi, Johnny Wei, Robin Jia. This research was supported in part by a Young Investigator award from the Allen Institute for AI.
@@
Appendix A Experimental Settings and Task Instructions
We compare three model pairs: GPT-3.5 vs. Vicuna-7B (LMSys, 2023), Vicuna-7B vs. Mistral-7B (Jiang et al., 2023), and GPT-3.5 vs. FLAN-T5-XXL (Longpre et al., 2023). [^10] Available via HuggingFace ModelHub: lmsys/vicuna-7b-v1.5, mistralai/Mistral-7B-Instruct-v0.2, and google/flan-t5-xxl. We experiment with a set of instruction-tuned LLMs that can be zero-shot prompted for the chosen tasks. We use identical instructions for each model, with different model-specific system prompts that were used to fine-tune each model during instruction tuning. We prompt each model in a zero-shot manner. These instruction prompts are listed in Appendix A. For our experiments, we use temperature sampling with temperature $\tau=0.5$. To calculate alignment scores, we use $K=5$ samples and $C=25$ cross-alignment comparisons, unless mentioned otherwise. In Table 1, we present the instruction prompts we used for each dataset described in §2.3.
For auto-evaluation experiments, we use: alpaca_eval_gpt4, alpaca_eval_cot_gpt4 _turbo_fn, and alpaca_eval_llama3_70b_fn from AlpacaEval (Dubois et al., 2024).
Similarity Functions
[…]
@@ Appendix A Experimental Settings and Task Instructions
Appendix B Human Study Details
We hired a pool of 30 raters (workers) from Amazon Mechanical Turk, all of whom were native English speakers. Each rater was hired based on participation in a qualification study. The raters were paid at a rate of \$1.20 per HIT, which was equal to roughly \$18 per hour using internally calculated time estimates for a single HIT. The order in which models’ summaries were shown in each pair was randomized in order to prevent positional bias. We present the HIT interface shown to AWS MTurk workers in Figure 10.
@@ Appendix E Using other similarity metrics
In Appendix A, we describe how different similarity functions can be used for different tasks, as well as to measure different dimensions of separability.
We present separability distributions in Figures 13 to 16 for ROUGE-1 (Lin, 2004), the original BERTScore (Zhang et al., 2019), entity similarity [^14] We calculate entity similarity between two generations by using spacy to extract named entities and taking the Jaccard Similarity of the set of entities from each generation, and cosine similarity [^15] Using sentence-transformers/all-distilroberta-v1 on HuggingFace ModelHub as our text embedding model., and text embedding cosine similarity on the summarization benchmarks used in our experiments.

### v3

@@ 1 Introduction
Finally, as a direct application of separability, we extend it to ELO, a rating system based on pairwise comparisons, now used widely for LLM generations (Chiang et al., 2024). By modifying the ELO rating update function to account for the separability of each new instance, our separability-ELO ratings provide more nuanced comparisons (§4). Overall, separability offers a reliable signal in the noisy landscape of generative evaluation via pairwise ratings, provides insights into benchmarks and test instances, and complements existing ranking measures for robust preference evaluation. Our code and data is available at https://github.com/dill-lab/separability.
@@
At a very high level, there are four common scenarios which may occur in comparing generations from two models, which we highlight in Figure 2.111While these scenarios aren’t fully exhaustive, they represent 97.5 % of the roughly 30,000 instances that are used for experiments in this section. Scenarios 1, 2 and 3 all depict output sets where any sample from model $A$ is expected to be very distant from any sample from model $B$ (i.e. low cross-alignment). It is easy to distinguish the two models under scenarios 1 and 3. In Scenario 1, if two generations are very different, they must be from different models. For Scenario 3, the difference in self-alignments is a clear distinguishing factor between the two sets of generations. Under scenario 2, generations from the same model are also far from each other (i.e. the self-alignment is also low), which makes the overall output sets hard to distinguish. In contrast, scenario 4 depicts a situation where both self- and cross-alignments are high; all generations, regardless of the model they came from, are similar, making it hard for the rater to distinguish the models’ output sets.
@@ 2.1 Calculating Generation Alignments
where $s:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ is a text similarity metric such as ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), or BLEU (Papineni et al., 2002). Intuitively, the alignment score $\mathcal{A}_{A,B}^{i}$ measures the expected similarity between an output from $m_{A}$ and an output from $m_{B}$, parameterized by $s$. A high value of $\mathcal{A}_{A,B}^{i}$ indicates high similarity (low variability) between the generations of the two models. Different similarity metrics can be used for different tasks. In cases where a user cares about fine-grained lexical differences, a metric such as BLEU may be suitable. On the other hand, if fine-grained lexical differences are less important than coarser semantic differences, metrics such as BERTScore or word embedding cosine similarity would be more suitable. We use a variation of BERTScore with a length-adjustment (defined in Section 2.2), unless otherwise noted.
@@ 2.2 Calculating Generation separability
separability can take values in $[-1,1]$. [^4] We apply min-max normalization to constrain the range of alignments to $[0,1]$ and to make alignment scores more interpretable and comparable across different model classes. In practice, however, cross-alignment is usually lower than self-alignment, making $\delta_{A,B}^{i}\in[0,1]$.
Similarity Functions
[…]
Figure 3: Histograms of separability distributions for summarization (Left) and translation (Right). For similar model pairs, CNN/DailyMail for news summarization and translation from a high-resource language (German) have lower average separability compared to SAMSum for dialogue summarization and translation from a lower-resource language (Czech). We use length-adjusted BERTScore Zhang et al. (2019) (defined in Section 2.2) as the similarity metric for summarization and BLEUPapineni et al. (2002) for translation.
Figure 4: separability distributions for ART and BiSECT. We use length-adjusted BERTScore here (defined in Section 2.2) as the similarity metric. separability has higher variance, especially for BiSECT, largely caused by differences in instruction prompt interpretation; see Appendix C.
@@ 2.3 Computing separability on Generation Benchmarks
We compute separability for various generation tasks under 6 different benchmarks using 3 model pairs, and demonstrate how separability can allow model developers and users to visualize and understand how much a model pair’s generations differ on a particular dataset. Figure 3 shows the empirical separability distributions on two summarization benchmarks (left): CNN/DailyMail (Nallapati et al., 2016) and SAMSum (Gliwa et al., 2019), and two machine translation benchmarks (right): Czech to English and German to English from the WMT-19 dataset (Barrault et al., 2019). Figure 4 shows the empirical separability distributions for abductive reasoning and sentence simplification, where we use the ART (Bhagavatula et al., 2019) and BiSECT (Kim et al., 2021) benchmarks respectively.
For each benchmark, we compare three model pairs: GPT-3.5 vs. Vicuna-7B (LMSys, 2023), Vicuna-7B vs. Mistral-7B (Jiang et al., 2023), and GPT-3.5 vs. FLAN-T5-XXL (Longpre et al., 2023). We use identical instructions for each model, with different model-specific system prompts that were used to fine-tune each model during instruction tuning. We prompt each model in a zero-shot manner. In Appendix A, we present the instruction prompts we used for each dataset described in this section. For our experiments, we use temperature sampling with temperature $\tau=0.5$. To calculate alignment scores, we use $K=5$ samples and $C=25$ cross-alignment comparisons, unless mentioned otherwise. In Appendix C, we present examples of low and high separability generations corresponding to each of these tasks.
@@
We conduct a human study to verify our formulation for separability as a meta-evaluation measure of preference rating consistency for a given instance $\mathbf{x}^{i}$ and a pair of generative models, $m_{A}$ and $m_{B}$. Given that generation sets corresponding to low separability instances are harder to distinguish, we hypothesize that preference judgments from raters on those sets will be inconsistent. In other words, raters will not consistently prefer the same model’s generation for any pair of generations sampled from low separability instances. On the other hand, we hypothesize that a rater’s preference judgments on generation sets corresponding to high separability instances will be consistent. While the inherent subjectivity present in generative evaluation may prevent raters from agreeing with each other on high separability instances, we hypothesize that each individual rater’s judgments will consistently favor one model’s generations.
Figure 6: Proportion of rating sets with each value in the range of consistency corresponding to different separability ranges. Ratings are not aggregated across raters for a test instance here. For each model pair and dataset configuration, the support is divided into equal sized ranges. The proportion of perfectly consistent ratings increases, and the proportion of inconsistent ratings decreases in higher separability ranges.
@@ 3.1 Rating Consistency
We define consistency of preference judgments as the average ratings from raters over $N$ sampled pairs. For an input instance $\mathbf{x}^{i}$, we sample $N$ generations $\mathbf{y}^{i,j}\in\mathcal{Y},j\in 1\ldots N$ each from models $A$ and $B$ to obtain a set of paired generations $\mathcal{P}_{AB}(\mathbf{x}^{i})=\left\{\left(\mathbf{y}^{i,j}_{A},\mathbf{y}^{i,j}_{B}\right)\right\}_{j=1}^{N}$.
@@ 3.2 Study Protocol and Settings
We conducted a human study with raters hired from Amazon Mechanical Turk. Each human intelligence task (HIT) consisted of reading a source text (in our case, a news article or a dialogue) and $N=5$ pairs of generated summaries. [^8] While we performed our experiments with summarization, we expect our results to hold for other tasks as well. In addition, the raters are performing $5$ cross-model comparisons as opposed to $25$ when calculating separability, but we find that $5$ comparisons suffice. For each summary pair, raters were asked to select which summary they preferred, with the option of picking no preference.
We hired a pool of 30 raters (workers) from Amazon Mechanical Turk, all of whom were native English speakers. Each rater was hired based on participation in a qualification study. The raters were paid at a rate of \$1.20 per HIT, which was equal to roughly \$18 per hour using internally calculated time estimates for a single HIT. The order in which models’ summaries were shown in each pair was randomized in order to prevent positional bias. The HIT interface can be found in Appendix B.
@@
Figure 8: Mean consistency of human and auto-rater preference judgments increases with separability. Mean consistency is computed over all $600$ HITs collected. Consistency for a particular test instance is aggregated over raters by taking the mean of each individual rater’s rating consistency.
@@ 3.4 Higher separability Instances Receive Consistent Auto-Rater Ratings
LLM-based automatic raters have been rising in popularity (Chang et al., 2024) and are being used to replace human raters in many preference evaluation setups. We ask: do auto-raters produce similar patterns of consistency as humans when making preference judgments? We repeat our experiments using the same 600 instances in §3.3 with auto-raters provided by AlpacaEval (Dubois et al., 2024) Each test instance is judged by three auto-raters, which have the highest correlations with humans (as of June 2024). [^9] These models are: alpaca_eval_gpt4, alpaca_eval_cot_gpt4 _turbo_fn, and alpaca_eval_llama3_70b_fn from AlpacaEval (Dubois et al., 2024). Since these raters cannot give tie judgments, the only possible consistencies are 00 or $1$.
@@ 4 Applying separability to ELO
We compute ELO and separability-weighted ELO (sep-ELO; Equation 6) using data from our 600 human evaluation HITs as preference judgments (§3). To calculate these ratings, we sample one rating for each input from our pool of ratings. We compute confidence intervals with the bootstrap method with $100$ trials. Figure 9 shows that sep-ELO has narrower gaps in model ranking, suggesting that models are more similar under adjustments to consistency of judgments (or, separability). We acknowledge that our results are a pilot due to the limited number of ratings we could use in our computation. However, we expect sep-ELO can reveal reliable trends even when applied to larger sets of preference data such as LMSYS [^12] https://lmsys.org/. Further testing with larger sets of ratings can also help users more optimally tune the hyperparameters used in the separability-ELO score calculation. However, we note that hyperparameter selection involves a subjective judgment on how much a practitioner wishes to incorporate separability values into the score update.
@@
Beyond sep-ELO, separability values could be used for adversarially filtering test sets Bras et al. (2020). Not only would this lead to fine-grained comparisons between models, but could also lead to obtaining cost- and time-efficient human ratings. However, some caution is to be urged since such filtering may lead to biases Schwartz and Stanovsky (2022), since low separability instances can still contain valuable information. Instead, we recommend importance weighting instances by separability when sampling instances for human judgments, in a similar manner as it is used in ELO ratings. That is, we recommend that high separability instances be used for human preference judgment collection. Alternatively, a stratified sampling approach from different separability ranges could ensure a more robust preference evaluation scheme.
@@ Acknowledgments
We would like to thank several members of USC NLP for their valuable feedback, insights, and help in designing and testing our human evaluation: Jaspreet Ranjit, Risha Surana, Xinyue Cui, Yoonsoo Nam, Keyu He, Joseph Liu, Matthew Finlayson, Brihi Joshi, Johnny Wei, and Robin Jia. This research was supported in part by a Young Investigator award from the Allen Institute for AI, as well as the National Science Foundation under Grant No. IIS-2403436. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@
Appendix A Task Instructions
We list the zero-shot instruction prompts we used for experiments in Section 2.3 in Table 1.
@@ Appendix A Task Instructions
Appendix B Human Study Interface
We present the HIT interface shown to AWS MTurk workers in Figure 10.
@@ Appendix E Using other similarity metrics
In Section 2.2, we describe how different similarity functions can be used for different tasks, as well as to measure different dimensions of separability.
We present separability distributions in Figures 13 to 16 for ROUGE-1 (Lin, 2004), the original BERTScore (Zhang et al., 2019), entity similarity [^13] We calculate entity similarity between two generations by using spacy to extract named entities and taking the Jaccard Similarity of the set of entities from each generation, and cosine similarity [^14] Using all-distilroberta-v1 in the sentence-transformers library, and text embedding cosine similarity on the summarization benchmarks used in our experiments.